OpenStudy (moongazer):

The zeros of f(x) = 72x^2-9x are? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

72x(x-9)=0 x=0 n x=9 are the zeros

OpenStudy (anonymous):

zeros of a polynomail are the values of the variable for which f(x)=0

OpenStudy (moongazer):

based on my answer key the answer is 0 and 1/8 i just don't know how it become like that

OpenStudy (anonymous):

oh m wrong 9x(8x-1)=0 x=0, 8x-1=0 x=1/8

OpenStudy (moongazer):

what happened to 9x??

OpenStudy (anonymous):

9x=0 x=0 coz 9 is not equal to zero

OpenStudy (moongazer):

so it means you will get the value of 2 x's x=0/9 which is x=0 and 8x-1=0 which is x=1/8 am i correct?

OpenStudy (anonymous):

right :)

OpenStudy (moongazer):

ok thanks!

OpenStudy (moongazer):

you earned a medal ^_^